OpenStudy (anonymous):

f(x)=x^2+4x-12 determine the zeros

OpenStudy (anonymous):

f(x) = (x-2)*(x+6), x = 2 or -6

OpenStudy (anonymous):

ok i am so bad at factoring, how exactly did you do this

OpenStudy (anonymous):

if its probably nice, then it is two factors of 12 as the a's and b's of the form (x+a)*(x+b). Since it is an x^2 out front, it will be the two factors that add up to 4, -2+6 = 4

OpenStudy (anonymous):

oh?! ok well I have a few more a little more difficult can you help? I can start a new question

OpenStudy (anonymous):

i gtg, sorry